# EMF Mitigation via 5G and 6G Scheduling

### v1

@@ V Conclusion
In this work we propose a family of techniques to spread the EIRP budget in each slot and then to allocate power and resources to UEs upon MAC scheduling slot-by-slot operations. Differently from MAC scheduling EIRP control prior art, we dynamically curb the power per UE in a computationally feasible manner. This is achieved by minimizing dependencies among different “segments” where EIRP constraints are active and by deriving a proper water-filling solution to the resulting parallel problems.
Our proposal allows to drastically reduce the communications performance losses compared to the legacy approaches that dynamically adapt the resource allocation without any power adaptation. For example, in case of $25\%$ allowed EIRP, our best proposal loses only $6.5\%$ in cell throughput, in contrast to the $44.2\%$ reduction with the legacy approaches. The minimal computational overhead and the performance achieved in our extremely detailed simulation study confirm that our technique is a standout candidate to enforce EIRP constraints in slot-by-slot operations of 5G and future 6G MAC schedulers.
@@ Acknowledgements
The authors would like to thank Tedros Abdu for the great discussions and feedback during the evolution of this study.

### v2

@@ V Conclusion
In this work we propose a set of techniques for EMF exposure mitigation that control the base station EIRP on a slot-per-slot basis to adhere to a budget defined over multiple slots. This budget is produced by an outer-loop algorithm such as the one in [9]. The integration of these two techniques enables a base station to comply with the “actual maximum approach” defined in [3]. Our solution reduces the transmission power fairly across all UEs with relatively low computational complexity. Moreover, it is fully parallelizable across different segments, that partition the sector into azimuth/elevation angle clusters on which configurable EIRP thresholds.
Our proposal significant alleviates the impact of EMF mitigation constraints on user performance compared to legacy approaches that only adjust resource allocation without any power adaptation. For example, when the power reduction factor is set to $1/4$ (or $-6$ dB), our method reduces cell throughput at high loads by a mere $6.5\%$, in contrast to a $44.2\%$ reduction with the legacy approaches. The minimal computational overhead and the performance achieved in our detailed massive-MIMO simulation study confirm that our technique is a standout candidate for enforcing actual EIRP constraints in slot-by-slot operations of 5G and future 6G MAC schedulers.
@@ Acknowledgements
The authors thank Tedros Abdu and Alois Herzog for the discussions and feedback during the evolution of this study.